# Coordinated Motion Control of a Wire Arc Additive Manufacturing Robotic System for Multi-directional Building Parts

### v1

@@ Abstract
This work is motivated by a common problem concerning wire arc additive manufacturing (WAAM): manufacturing parts with variable build directions. The aim of this work is to provide, for both practitioners and researchers of the additive manufacturing field, a closed loop control algorithm for manufacturing parts with a dual robot system. One robot holds the torch, while the other has a substrate attached to its end. The proposed robots’ coordinated motion can perform the deposition while the workpiece is repositioned guaranteeing that the torch orientation is being kept aligned with a reference axis (e.g. gravity) given in an inertial frame.
Index Terms:
Wire Arc Additive Manufacturing, WAAM, Coordinated motion control, Task augmentation, Constrained Jacobian
@@ I Introduction
Although MAM cannot be considered a novel method of manufacturing parts, since the first experiments on this technology date from the early 1980s when it was first termed as rapid prototyping [2], it has recently attracted interest from industry and academia due to its high capability to manufacture complex structural parts that are difficult or even not possible to manufacture through conventional manufacturing techniques [2, 3].
@@
Among the available MAM technologies described in the ASTM F2792 standard [5], wire arc additive manufacturing (WAAM) stands out due to its potential capability to produce large parts with distinct geometry without the necessity of specific tools. WAAM is the combination of wire feedstock with an electric arc as a heat source, all guided by a robotic system to enable the buildup of a designed part. It uses standard robotic welding hardware, such as welding power source, torch, wire feeding system, positioning table, robot arm, and others common welding equipment [6], which makes it less expensive and more adaptable for existing production lines.
@@
In order to produce complex geometry parts and reduce production time, [14], [15] proposed similar offline planning algorithms for manufacturing specific rounded metal additive manufacturing parts parts with a system that coupled a 6-axis robot manipulator with a 2-axis positioning table. These methods consist of planning the deposition trajectory while using an external-axis system in order to maintain the molten pool orientation with respect to gravity.
For these curved parts offline solutions [14, 15, 12], the part was either fabricated in the center of the deposition surface, aligned with the axis of rotation of the positioning table, or the substrate (e.g. metal cylinder) was centered to facilitate the part manufacturing and trajectory planning.
A general limitation of offline planning methods is that although they allow parameters correction before each layer deposition, they cannot make trajectory or deposition parameters corrections during the deposition. For instance, the deposition bead/layer geometry could be monitored and estimated by sensors and an online closed loop control could adjust deposition parameters (arc current, wire feed speed), deposition path and tool center point (TCP) travel speed [16].
@@ III Differential kinematics model
where the TCP velocities vector $\bm{v}$ is composed by its linear $\bm{\dot{p}}\in\mathbb{R}^{3}$ and angular $\bm{\omega}\in\mathbb{R}^{3}$ velocities with respect to $\mathcal{F}_{d}$ represented in $\mathcal{F}_{d}$. This relationship can also be expressed by partitioning the Jacobian matrix $J(\bm{\theta})$ into column blocks corresponding to each robot joints:
@@ IV-A Augmented tasks for WAAM redundant systems
The Jacobian $J_{A}(\bm{\theta})$ is bounded for all $\bm{\theta}$, that is, $||J_{A}(\bm{\theta})||_{\infty}\leq c_{0}$, $\forall{\theta}_{i}\in[0,2\pi)$, where $c_{0}>0$ and $i=1,\dots,m$.
@@ IV-B Control law
The system (21) is valid if and only if, the following assumptions hold: (A1) the robot kinematics is perfectly known; (A2) the control law $[\bar{\bm{u}}_{1}^{\mathsf{T}}\quad\bar{\bm{u}}_{2}^{\mathsf{T}}]^{\mathsf{T}}$ does not drive the robotic system to singular configurations.
@@
Now considering $\bm{z}^{b}$, $z$-axis of $\mathcal{F}_{t}$, and the desired $\bm{z}^{b}_{d}$ direction both represented in $\mathcal{F}_{t}$, the orientation error is given by
@@
[…]
@@
[…]
@@ V-B Equivalence with the Augmented Jacobian
where $J_{A}\!\in\!\mathbb{R}^{6+r\times m}$ is the same augmented Jacobian (17) that maps the joint velocities signal to the augmented velocity vector $[\bm{v}^{\mathsf{T}}\bm{\omega}_{s}^{\mathsf{T}}]^{\mathsf{T}}$.
@@ VI Singularity Analysis
For the WAAM system of Fig. 3 and augmented task problem defined in (17), this algorithmic singularity arise when $\mathcal{F}_{d}$ has its $z$-axis aligned with the inertial frame used to design the secondary task. Algebraically that means that $J$ and $\Lambda J_{2}$ have linearly dependents rows and the augmented Jacobian $J_{A}$ is not invertible. This becomes obvious in the constrained Jacobian algorithm, since this configuration leads the first column of $J_{1}$ to become linearly depend of the columns of $\Lambda^{\#}$.
@@
The singularity occurs when $k<n+r$ and the Jacobian has null singular values. A common algorithm to reduce the effects of the inversion of the ill-contioned Jacobian is the Damped Least Square (DLS) method [28] where the control law (21) is rewritten as:
@@ VII-C Curved thin-wall
where $n_{l}$, $l_{h}$, $l_{l}$ , $o_{x}$ and $o_{y}$ follows the description of (44), and $o_{z}$ is the offset along $\mathcal{F}_{d}$ $z$-axis (chosen to be the first section layers average height). The incremental inclination angle $\Delta\gamma$ is computed according to the curve radius $R$ and the layer height $l_{h}$ as:
@@
A curved wall part manufactured with this method is show in Fig. 7(b). The final part has a total length of $l_{l}\approx 97$ mm, divided into segments where $h_{1}=30$ mm, $R=30$ mm and $h_{2}=20$ mm. The travel speed used for this experiment is set at $V_{ts}=5$ mm/s with an acceleration and deceleration time of $t_{a}=0.1$ s. The deposition parameters produce beads with an average height of $2.0$ mm, resulting in a final part composed of approximately $n_{t}\approx 48$ layers.
@@
It is worth to point out that, during the curved thin wall manufacturing, some adjustments were made regarding the torch orientation with respect to gravity. This was due to the fact that the torch could not satisfy the desired orientation without colliding with some of the screws holding the substrate to the base plate. Although this issue was unintentional, the proposed framework allowed for an easy adjustment of the torch direction relative to gravity, highlighting another advantage of the method. To address this issue, the torch orientation with respect to $\mathcal{F}_{ab}$ has been changed to rotate $\pi/18$ ($10^{\circ}$) around $y$ axis, resulting in $z_{d}\approx[-0.17\quad 0\quad-0.98]^{\mathsf{T}}$.
@@ VII-D Intake funnel
The travel speed of $V_{ts}=5$ mm/s produces layers with average height of $2.0$ mm. The planned part consists of a cylinder with $r_{c}=80$ mm radius and $h\approx 90$ mm height (which requires the deposition of $45$ layers). For the second section has a $\pi/3$ ($60^{\circ}$) opening flare with a radius of $r=20$ mm, requiring the deposition of $21$ layers as each layer height is planned to have $l_{h}=2$ mm.
@@ VIII Concluding Remarks and Perspectives
Future works include expressing the coordination control for WAAM as an optimization problem, using slack variables to ease the trajectory tracking of some degrees of freedom allowing the obstacle avoidance for depositions in non-regular substrate or cladding of cylindrical parts. The authors also intend to investigate the use of the strategy for higher redundancy levels systems aimed at hybrid manufacturing applications. Proof of theorem 1 Consider the Lyapunov candidate function

### v2

@@ Abstract
This work addresses the technical constraints of Wire Arc Additive Manufacturing (WAAM) regarding non-planar geometries and variable build directions. To mitigate issues like unsupported overhangs and the staircase effect, we present a closed-loop coordinated control algorithm for dual-robot systems, where a manipulator guides the torch and a secondary robot repositions the substrate. The practical advantage lies in the task-augmentation scheme that automates gravity-aligned deposition, ensuring the melt pool remains stable and flat throughout the process. This eliminates the need for complex support structures and minimizes post-processing. By synchronizing both robots, the algorithm allows operators to plan trajectories directly on the workpiece frame while the controller autonomously maintains optimal torch orientation relative to an inertial frame, significantly enhancing surface quality and operational efficiency in complex part production.
@@ I Introduction
Although MAM cannot be considered a novel method for manufacturing parts, since the first experiments with this technology date from the early 1980s, when it was first termed rapid prototyping [2], it has recently attracted interest from industry and academia due to its high capability to manufacture complex structural parts that are difficult or even not possible to manufacture using conventional manufacturing techniques [2, 3].
@@
Among the available MAM technologies described in the ASTM F2792 standard, wire arc additive manufacturing (WAAM) stands out due to its potential capability to produce large parts with distinct geometry without the necessity of specific tools. WAAM is the combination of wire feedstock with an electric arc as a heat source, all guided by a robotic system to enable the buildup of a designed part. It uses standard robotic welding hardware, such as welding power source, torch, wire feeding system, positioning table, robot arm, and others common welding equipment [5], which makes it less expensive and more adaptable for existing production lines.
@@
In order to produce complex geometry parts and reduce production time, [13] proposed similar offline planning algorithms for manufacturing specific rounded metal additive manufacturing parts parts with a system that coupled a 6-axis robot manipulator with a 2-axis positioning table. These methods consist of planning the deposition trajectory while using an external-axis system in order to maintain the molten pool orientation with respect to gravity.
For curved parts offline solutions [14, 13, 11], the part was either fabricated in the center of the deposition surface, aligned with the axis of rotation of the positioning table, or the substrate (e.g. metal cylinder) was centered to facilitate the part manufacturing and trajectory planning.
A general limitation of offline planning methods is that although they allow parameters correction before each layer deposition, they cannot make trajectory or deposition parameters corrections during the deposition. For instance, the deposition bead/layer geometry could be monitored and estimated by sensors and an online closed loop control could adjust the deposition parameters (arc current, wire feed speed), deposition path and tool center point (TCP) travel speed [15, 16].
@@ III Differential kinematics model
where the TCP velocities vector $\bm{v}$ is composed of its linear $\bm{\dot{p}}\in\mathbb{R}^{3}$ and angular $\bm{\omega}\in\mathbb{R}^{3}$ velocities with respect to $\mathcal{F}_{d}$ represented in $\mathcal{F}_{d}$. This relationship can also be expressed by partitioning the Jacobian matrix $J(\bm{\theta})$ into column blocks corresponding to each robot joints:
@@ IV-A Augmented tasks for WAAM redundant systems
The Jacobian $J_{A}(\bm{\theta})$ is bounded for all $\bm{\theta}$, that is, $J_{A}(\bm{\theta})||_{\infty}\leq c_{0}$, $\forall{\theta}_{i}\in[0,2\pi)$, where $c_{0}>0$ and $i=1,\dots,m$.
@@ IV-B Control law
The system (21) is valid if and only if, the following assumptions hold: (A1) the robot kinematics is perfectly known; (A2) the control law $\bar{\bm{u}}_{1}^{\mathsf{T}}\quad\bar{\bm{u}}_{2}^{\mathsf{T}}]^{\mathsf{T}}$ does not drive the robotic system to singular configurations.
@@
Now considering $\bm{z}^{b}$, $z$-axis of $\mathcal{F}_{t}$, and the desired $\bm{z}^{b}_{d}$ direction both represented in $\mathcal{F}_{t}$, the orientation error is given by:
@@
where $\alpha=\cos^{-1}(\bm{z}^{b}_{d}\cdot\bm{z}^{b})$ and $\bm{r}=(\bm{z}^{b}\times\bm{z}^{b}_{d})/\lVert\bm{z}^{b}\times\bm{z}^{b}_{d}\rVert$ are the angle and axis of the equivalent rotation needed to align $\bm{z}^{b}$ and $\bm{z}^{b}_{d}$. For the special case of aligning the TCP $z$-axis with a given reference with respect to $\mathcal{F}_{ab}$, $\bm{r}=[r_{1}\ r_{2}\ 0]^{\mathsf{T}}$ and $\bm{e}_{\bm{\epsilon}s}=[e_{\epsilon_{1}s}\ e_{\epsilon_{2}s}\ 0]^{\mathsf{T}}=[\bm{e}_{s}\ 0]^{\mathsf{T}}$.
@@
Considering the property P1, $J_{A}(\bm{\theta})\!\in\!\cal{L}_{\infty}$, and given that $\bm{\eta}\!\in\!\mathcal{L}_{2}\cap\cal{L}_{\infty}$, then $J_{A}(\bm{\theta})\bm{\eta}\!\in\!\mathcal{L}_{2}\cap\cal{L}_{\infty}$ [24]. Thus, it can be considered that there exists a positive constant $L_{m}$ such that $\int_{0}^{t}\bm{\eta}^{\mathsf{T}}J_{A}^{\mathsf{T}}J_{A}\bm{\eta}d\tau\leq L_{m}~\forall~t$. Therefore, the following theorem with properties of the closed loop system stability is presented:
@@ V-B Equivalence with the Augmented Jacobian
where $J_{A}\!\in\!\mathbb{R}^{6+r\times m}$ is the same augmented Jacobian (17) that maps the joint velocities signal to the augmented velocity vector $\bm{v}^{\mathsf{T}}\bm{\omega}_{s}^{\mathsf{T}}]^{\mathsf{T}}$.
@@ VI Singularity Analysis
For the WAAM system of Fig. 3 and the augmented task problem defined in (17), this algorithmic singularity arises when $\mathcal{F}_{d}$ has its $z$-axis aligned with the inertial frame used to design the secondary task. Algebraically that means that $J$ and $\Lambda J_{2}$ have linearly dependent rows and the augmented Jacobian $J_{A}$ is not invertible. This becomes obvious in the constrained Jacobian algorithm, since this configuration leads the first column of $J_{1}$ to become linearly dependent of the columns of $\Lambda^{\#}$.
@@
The singularity occurs when $k<n+r$ and the Jacobian has null singular values. A common algorithm to reduce the effects of the inversion of the ill-conditioned Jacobian is the Damped Least Square (DLS) method [28] where the control law (21) is rewritten as:
@@ VII-C Curved thin-wall
where $n_{l}$, $_{h}, ll$$_{l}$ , $o_{x}$ and $o_{y}$ follows the description of (44), and $o_{z}$ is the offset along $\mathcal{F}_{d}$ $z$-axis (chosen to be the first section layers average height). The incremental inclination angle $\Delta\gamma$ is computed according to the curve radius $R$ and the layer height $l_{h}$ as:
@@
A curved wall part manufactured with this method is show in Fig. 7(b). The final part has a total length of $l_{l}\approx 97$ mm, divided into segments where $h_{1}=30$ $m$, $R=30$ mm and $h_{2}=20$ mm. The travel speed used for this experiment is set at $V_{ts}=5$ mm/s with an acceleration and deceleration time of $t_{a}=0.1$ s. The deposition parameters produce beads with an average height of $2.0$ mm, resulting in a final part composed of approximately $n_{t}\approx 48$ layers.
@@
It is worth to point out that, during the curved thin wall manufacturing, some adjustments were made regarding the torch orientation with respect to gravity. This was due to the fact that the torch could not satisfy the desired orientation without colliding with some of the screws holding the substrate to the base plate.
Although this issue was unintentional, the proposed framework allowed for an easy adjustment of the torch direction relative to gravity, highlighting another advantage of the method. To address this issue, the torch orientation with respect to $\mathcal{F}_{ab}$ has been changed to rotate $\pi/18$ ($10^{\circ}$) around $y$ axis, resulting in $z_{d}\approx[-0.17\quad 0\quad-0.98]^{\mathsf{T}}$.
@@ VII-D Intake funnel
The travel speed of $V_{ts}=5$ mm/s produces layers with average height of $2.0$ mm. The planned part consists of a cylinder with $r_{c}=80$ mm radius and $h\approx 90$ mm height (which requires the deposition of $45$ layers). For the second section has a $\pi/3$ ($60^{\circ}$) opening flare with a radius of $r=20$ mm, requiring the deposition of $21$ layers as each layer height is planned to have $l_{h}=2$ $m$.
@@ VIII Concluding Remarks and Perspectives
Future works include expressing the coordination control for WAAM as an optimization problem, using slack variables to ease the trajectory tracking of some degrees of freedom allowing the obstacle avoidance for depositions in non-regular substrate or cladding of cylindrical parts. The authors also intend to investigate the use of the strategy for higher redundancy levels systems aimed at hybrid manufacturing applications.
Proof of theorem 1 Consider the Lyapunov candidate function